OpenStudy (lilkg77):

PLEASE HELP!!!! which of these points lies on the line described by the equation below? y-5=6(x-7) A. (5, 7) B. (7, 5) C. (-7, -5) D. (-5, -7)

pooja195 (pooja195):

\[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

pooja195 (pooja195):

x1 is located where 7 is y1 is located where 5 is what are the coordinates?

OpenStudy (lilkg77):

7, 5?

pooja195 (pooja195):

Correct :)

OpenStudy (lilkg77):

Thanks :)

pooja195 (pooja195):

np =)